OpenStudy (anonymous):

How would i solve this please. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

20+25+50x-100=145 parenthesis first +100=+100 add 100 to both sides 20+25+50x=245 45+50x=245 -45 -45 subtract 45 from each side 50x=200 --- --- divide by 50 50 50 x=4